Shift report 20011118 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Nov 18 2001 - 08:16:00 EST

  • Next message: Fouad Rami: "Re: run status"

    ------------------------------------------------------------
    Date: Sun Nov 18 08:16:00 2001, Shift: 00:00-08:00
    Supervisor: Y.K.Lee, with: 
    Log book#11 pages 7 - 7
    
    Shift Summary:
    ------------------------------------------------------------
    1. Good beam till 3 AM. No more beam for the rest
    of the shift due to RF problem at AGS.
    
    2. Continue data with MRS @90 with 6B, and FS @12 with 2/3A.
       Run 5868 has 165K trigger 6.
       Run 5869 has 63K trigger 6.
       Runs 5870,5871,5872 are Detectors pedestal, TPC pedestal
          and DVM pedestal runs, respectively.
    
    3. No new detector problems.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Nov 18 2001 - 08:16:23 EST